at 9:00 on Saturday two bicyclist heading in opposite directions pass each other on a bicycle path. the bicyclist heading north is riding 6km/hr faster than the bicyclist heading south. at 10:15,they are 42.5km apart. find the two bicyclists rates.
Answer by mananth(16949)

at 9:00 two bicyclist heading in opposite directions
the bicyclist heading south x km/h
bicyclist heading north x+6 km/hr.
at 10:15,= 1 1/4 hour = 5/4 hours
they are 42.5km apart.
.....
5/4 *x +5/4 *(x+6) = 42.5
5x/4+(5x+30) /4 =42.5
(5x+5x+30)=4*42.5
10x+30=170
10x=170-30
10x=140
x=14 km/h south
14 +6 = 20km/h north bound